The amount of time it takes p people to paint d doors varies directly with the number of doors and inversely with the number of people. Four people can paint 10 doors in 2 hours. How many people will it take to paint 25 doors in 5 hours?
Answer by rothauserc(4718)

time = k * ( d / p ) where k is the constant of proportionality
****************************************************************
2 = k * (10/4)
10k = 8
k = 8/10 = 4/5
****************************************************************
5 = (4/5) * (25/p)
5 = 100 / 5p
25p = 100
p = 4
therefore
4 people will paint 25 doors in 5 hours